QUESTION 1

The sample size is not one of the causes of committing an statistical error.

True

False

QUESTION 2

If the alternative hypothesis has a mathematical sign of above (>) and the experimental Z value is +2.52 then the p – value is 0.994132.

True

False
Answer #1

(1) False

(2) False, because P-value=P(z>2)=1-P(z<=2) = 0.0228